# FLOAT: Generative Motion Latent Flow Matching for Audio-driven Talking Portrait

### v4

@@ Abstract
$\star$\$$\star$\$footnotetext: This work was done during South Korea Mandatory Military Service at DeepBrain AI Inc.
@@ 3 Preliminaries: (Conditional) Flow Matching
Figure 2: Overview of FLOAT. We encode the source image $S\in\mathbb{R}^{3\times H\times W}$ into the latent with the explicit identity-motion decomposition $w_{s}=w_{s\to r}+w_{r\to s}\in\mathbb{R}^{d}$. Given audio segments $a^{-L^{\prime}:L}\in\mathbb{R}^{(L^{\prime}+L)\times d_{a}}$ of the length $L^{\prime}+L$ and the reference motion $w_{r\to s}$in $\mathbb{R}^{d}$, and the speech-driven emotion label $w_{e}$in $\mathbb{R}^{7}$, a flow matching transformer estimates the generating vector field $v_{t}(\varphi_{t}(x_{0}),\mathbf{c}_{t};\theta)\in\mathbb{R}^{L\times d}$ from noisy motion latents, which is used to solve corresponding ODE and generates the motion latents $w_{r\to\hat{D}^{1:L}}$. Finally, the sequence of latents $w_{S\to\hat{D}^{1:L}}:=(w_{S\to r}+w_{r\to\hat{D}^{l}})_{l=1}^{L}$ are decoded into the video $\hat{D}^{1:L}\in\mathbb{R}^{L\times 3\times H\times W}$.
Let $x$in $\mathbb{R}^{d}$ be a data, $t$in [0,1][0,1][ 0 , 1 ] be the time, and $q$ be a unknown target distribution. We can define a flow as a time-dependent transformation $\varphi_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that transforms a tractable prior distribution $p_{0}$ to the distribution $p_{1}\approx q$. This flow $\varphi_{t}$ further introduces a probability flow path $p_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{>0}$ and a generating vector field $v_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ where $p_{t}$ is defined by the push-forwarding
@@
where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x)$. However, the target generating vector field $u_{t}$ and the sample distribution $p_{t}$ are intractable. To address this issue, [42] proposes a method for constructing a “conditional" probability path $p_{t}(\cdot|x_{1})$ as well as target “conditional" vector field $u_{t}(\cdot|x_{1})$ using a sample $x_{1}$sim $q$ as a condition. And they prove that the following objective
@@
where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x|x_{1})$, is equivalent to (3) with respect to the gradient $\nabla_{\theta}$.
One natural way of constructing $u_{t}(\cdot|x_{1})$ is a “straight line" that connects $x_{0}$sim $p_{0}$ and $x_{1}$sim $q$, drawing an optimal transport (OT) path with constant velocity [42]. Specifically, a linear time interpolation between $x_{0}$ and $x_{1}$ gives us the flow $x_{t}=\varphi_{t}(x)=(1-t)x_{0}+tx_{1}$, the conditional probability path $p_{t}(x|x_{1})$ defined via the affine transformation $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-t)^{2}I)$, and the target generating vector field $u_{t}(x|x_{1})=x_{1}-x_{0}$. This specific choice turns the objective (4) into
@@
where $t$sim $\mathcal{U}[0,1]$, $x_{0}$sim $p_{0}$, and $x_{1}$sim $q$, all of which are tractable.
@@ 4 Method: Flow Matching for Audio-driven Talking Portrait
We provide an overview of FLOAT in Fig.˜2. Given source image $S$in $\mathbb{R}^{3\times H\times W}$, and a driving audio signal $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of length $L$, our method generates a video
@@ 4.1 Motion Latent Auto-encoder
Instead of VAE of SD, we pre-train a motion latent auto-encoder that encodes images into motion latents. This is done by training an auto-encoder to reconstruct a driving image from a source image sampled from the same video clip, enforcing the encoder to implicitly capture both temporally adjacent and distant motions. Inspired by LIA [85], we introduce a learned orthonormal basis that decomposes the motion along distinct orthogonal directions. Specifically, our motion auto-encoder encodes the source $S$ into the latent $w_{S}\in\mathbb{R}^{d}$ with following explicit decomposition:
@@
where $w_{S\to r}$in $\mathbb{R}^{d}$ is the identity latent and
@@
is the motion latent with $\lambda(S)$ :=:=:= $\left(\lambda_{m}(S)\right)_{m=1}^{M}$in $\mathbb{R}^{M}$ being the source-dependent motion coefficients that span the learned source-agnostic motion basis $V:=\{\mathbf{v}_{m}\}_{m=1}^{M}\subseteq\mathbb{R}^{d}$. In this space, $\lambda_{m}(S)$ is the intensity of the motion direction $\mathbf{v}_{m}$.
As shown in Fig.˜6, our method enables motion editing of the sampled (generated) motion using only the basis $V$ and its orthogonality, as stated in Eq.˜15.
@@ 4.2 Flow Matching in Motion Latent Space
Armed with this linear orthogonal space, we employ OT-based flow matching [42, 44] for the motion sampling. Specifically, we predict a vector field $v_{t}(x_{t},\mathbf{c}_{t};\theta)$in $\mathbb{R}^{L\times d}$ where $x_{t}$ is the sample at flow time $t\in[0,1]$, and $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ represents the driving conditions for $L$ consequent frames. This vector field generates the flow $\varphi_{t}:[0,1]\times\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ of $L$ frames by solving ODE (Eq.˜2). As illustrated in Fig.˜4, we build our vector field predictor upon the transformer encoder [79] architecture. Specifically, we adopt DiT [55] architecture, but decouple frame-wise conditioning from time-axis attention mechanism, which enables us to model temporally consistent motion latents.
@@
[…]
@@
Driving Conditions   We concatenate the audio representation $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of a pre-trained Wav2Vec2.0 [1], the speech emotion label $w_{e}$in $\mathbb{R}^{7}$, and the source motion latent $w_{r\to S}$in $\mathbb{R}^{d}$. Next, we add the flow time step embedding $\text{Emb}(t)$in $\mathbb{R}^{h}$ to these conditions, producing $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ via a linear layer, ToCondition, as depicted in Fig.˜2, where $\text{Emb}(t)$ is computed using the sinusoidal position embedding [79].
Training   We train FLOAT by reconstructing a target vector field computed from driving frames using the corresponding audio segments and a source motion latent. We choose a pair of driving motions and corresponding audio ($w_{r\to D^{1:L}}$, $a^{1:L}$), and construct the target vector field $u_{t}(x|w_{r\to D^{1:L}})=w_{r\to D^{1:L}}-x_{0}$in $\mathbb{R}^{L\times d}$ with noisy input $\varphi_{t}(x_{0})=(1-t)x_{0}+tw_{r\to D^{1:L}}$ ($t$sim $\mathcal{U}[0,1]$ and $x_{0}$sim $\mathcal{N}(0^{1:L},I)$).
@@
where $x_{t}$ :=:=:= $[w_{r\to D^{-L^{\prime}:0}}|~\varphi_{t}(x_{0})]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ is the concatenated input, $\mathbf{c}_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times h}$ is the driving condition consisting of $[t,w_{r\to S},w_{e},a^{1:L},a^{-L^{\prime}:0}]$. Note that $w_{e}$ and $w_{r\to S}$ are shared across the $L^{\prime}+L$ frames. We incorporate a velocity loss [75] to supervise temporal consistency:
@@
where $\Delta v_{t}$ and $\Delta u_{t}$ are the one-frame difference along the time-axis for the prediction $v_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ and the target $[w_{r\to D^{-L^{\prime}:0}}|~u_{t}]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$, respectively.
@@
where $\lambda_{\text{OT}}$ and $\lambda_{\text{vel}}$ are the balancing coefficients. During training, we apply dropout to $w_{r}$, $w_{e}$, and $a^{1:L}$ with a probability of 0.10.10.1 for CFV. Additionally, we apply dropout to the preceding audio and motion latents with a probability 0.50.50.5 for smooth transition in the initial window.
Table 1: Quantitative comparison results with state-of-the-art methods on HDTF [97] / RAVDESS [46]. The best result for each metric is in bold, and the second-best result is underlined. : evaluated with raw $\times 256$ 256 resolution outputs.
@@ 5.2 Implementation Details
[…]
@@ 5.4 Applications
Test-time Pose Editing via Orthonormal Basis $V$   Since FLOAT learns the underlying motion latent structure, it is natural to assume that for any sampled motion latent $w_{r\to\hat{D}}$, there exist motion coefficients $\{\lambda_{m}(\hat{D})\}_{m=1}^{M}$ satisfying the representation in Eq.˜9: $w_{r\to\hat{D}}$ $=$ $\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m}$.
@@
$\langle w_{r\to\hat{D}},~\mathbf{v}_{k}\rangle=\langle\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m},~\mathbf{v}_{k}\rangle=\lambda_{k}(\hat{D}),$ (15)
@@
Additional Driving Signals   In Fig.˜7 and Tab.˜2, we experiment with additional driving conditions, head poses and image-driven emotion labels, to explore additional controllability in our method. We employ 3DMM head pose parameters $p$in $\mathbb{R}^{6}$ [2] extracted by [13]. We concatenate a sequence of pose parameters $p^{1:L}$in $\mathbb{R}^{L\times 6}$ with the other driving conditions, and then map them to $c_{t}^{1:L}$in $\mathbb{R}^{L\times h}$. We also experiment on image-driven emotion [63] for frame-wise emotion control rather than the long-term emotion enhancement. FLOAT can effectively accommodate these additional conditions, highlighting its flexibility across diverse control signals.
@@ A.1 Model
Table 4: Quantitative comparison result (Same-identity) of motion latent auto-encoders on HDTF [97] / RAVDESS [46] / VFHQ [88]. The best result for each metric is in bold. : Results generated by official implementation ($\times 256$ 256)
@@ A.3 Training Objective
$\mathcal{L}_{\text{full-adv}}=-\log[\text{Disc}_{\text{full}}(\hat{D})],$ (19)
@@ B.2 Baselines
  • It is worth noting that we compare with two superior works EMO [76] and VASA-1 [90] based on their demonstration videos due to their unavailable implementation. We highly recommend referring to ‘01_\__EMO_\__VASA-1_\__Comparison///xxxx.mp4’.
@@ B.3 More on Experiments
Ablation on NFE   In general, increasing the number of function evaluation (NFE) reduces the solution error of ODEs. As shown in Tab.˜5, even with small NFE =2=2= 2, FLOAT can achieve competitive image quality (FID) and lip synchronization (LSE-D). However, it struggles to capture consistent and expressive motions (FVD and E-FID), resulting in shaky head motion and a static expression. This is because FLOAT generates the motion in the latent space, while image fidelity is determined by the auto-encoder. We provide supplementary videos, illustrating the impact of different NFE (Number of Function Evaluations). Notably, with a small NFE of 2, the generated images exhibit good quality, but the head movements appear temporally unstable, and emotions may be exaggerated. Please refer to supplementary videos for temporal jitters of low NFE.
@@
In Fig.˜20, we visualize the effect of different emotion guidance scale $\gamma_{e}$. For this experiments, the predicted speech-to-emotion label is disgust with $\%9$ probability. Notably, as increasing $\gamma_{e}$ from 0 to 2, we can observe that emotion-related expressions and motions are enhanced.
@@
[…]
@@
where $t$sim $\mathcal{U}[0,1]$, $\epsilon$sim $\mathcal{N}(0^{-L^{\prime}:L},I)$, and the noise input $x_{t}$in $\mathbb{R}^{(L^{\prime}+L)\times d}$ is sampled from a forward diffusion process $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ [27]. In our case, $x_{t}$ is noisy motion latents at diffusion time step $t$, starting from $t=0$ with $x_{0}$ $=$ $w_{r\to D^{1:L}}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$.

### v5

@@ Abstract
[…]
@@ 3 Preliminaries: (Conditional) Flow Matching
Figure 2: Overview of FLOAT. We encode the source image $S\in\mathbb{R}^{3\times H\times W}$ into the latent with the explicit identity-motion decomposition $w_{s}=w_{s\to r}+w_{r\to s}\in\mathbb{R}^{d}$. Given audio segments $a^{-L^{\prime}:L}\in\mathbb{R}^{(L^{\prime}+L)\times d_{a}}$ of the length $L^{\prime}+L$ and the reference motion $w_{r\to s}$ $\in$ $\mathbb{R}^{d}$, and the speech-driven emotion label $w_{e}$ $\in$ $\mathbb{R}^{7}$, a flow matching transformer estimates the generating vector field $v_{t}(\varphi_{t}(x_{0}),\mathbf{c}_{t};\theta)\in\mathbb{R}^{L\times d}$ from noisy motion latents, which is used to solve corresponding ODE and generates the motion latents $w_{r\to\hat{D}^{1:L}}$. Finally, the sequence of latents $w_{S\to\hat{D}^{1:L}}:=(w_{S\to r}+w_{r\to\hat{D}^{l}})_{l=1}^{L}$ are decoded into the video $\hat{D}^{1:L}\in\mathbb{R}^{L\times 3\times H\times W}$.
Let $x$ $\in$ $\mathbb{R}^{d}$ be a data, $t$ $\in$ $[0,1]$ be the time, and $q$ be a unknown target distribution. We can define a flow as a time-dependent transformation $\varphi_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that transforms a tractable prior distribution $p_{0}$ to the distribution $p_{1}\approx q$. This flow $\varphi_{t}$ further introduces a probability flow path $p_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{>0}$ and a generating vector field $v_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ where $p_{t}$ is defined by the push-forwarding
@@
where $t$ $\sim$ $\mathcal{U}[0,1]$ and $x$ $\sim$ $p_{t}(x)$. However, the target generating vector field $u_{t}$ and the sample distribution $p_{t}$ are intractable. To address this issue, [42] proposes a method for constructing a “conditional" probability path $p_{t}(\cdot|x_{1})$ as well as target “conditional" vector field $u_{t}(\cdot|x_{1})$ using a sample $x_{1}$ $\sim$ $q$ as a condition. And they prove that the following objective
@@
where $t$ $\sim$ $\mathcal{U}[0,1]$ and $x$ $\sim$ $p_{t}(x|x_{1})$, is equivalent to (3) with respect to the gradient $\nabla_{\theta}$.
One natural way of constructing $u_{t}(\cdot|x_{1})$ is a “straight line" that connects $x_{0}$ $\sim$ $p_{0}$ and $x_{1}$ $\sim$ $q$, drawing an optimal transport (OT) path with constant velocity [42]. Specifically, a linear time interpolation between $x_{0}$ and $x_{1}$ gives us the flow $x_{t}=\varphi_{t}(x)=(1-t)x_{0}+tx_{1}$, the conditional probability path $p_{t}(x|x_{1})$ defined via the affine transformation $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-t)^{2}I)$, and the target generating vector field $u_{t}(x|x_{1})=x_{1}-x_{0}$. This specific choice turns the objective (4) into
@@
where $t$ $\sim$ $\mathcal{U}[0,1]$, $x_{0}$ $\sim$ $p_{0}$, and $x_{1}$ $\sim$ $q$, all of which are tractable.
@@ 4 Method: Flow Matching for Audio-driven Talking Portrait
We provide an overview of FLOAT in Fig.˜2. Given source image $S$ $\in$ $\mathbb{R}^{3\times H\times W}$, and a driving audio signal $a^{1:L}$ $\in$ $\mathbb{R}^{L\times d_{a}}$ of length $L$, our method generates a video
@@ 4.1 Motion Latent Auto-encoder
Instead of VAE of SD, we employ LIA [85] as a base motion latent auto-encoder and pre-train it to encode images into motion latents. This is achieved by training the auto-encoder to reconstruct a driving image from a source image sampled from the same video clip, enforcing the encoder to implicitly capture both temporally adjacent and distant motions. Following [85], we use a learned orthonormal basis that can decompose the motion along distinct orthogonal directions. Specifically, our motion auto-encoder encodes the source $S$ into the latent $w_{S}\in\mathbb{R}^{d}$ with following explicit decomposition:
@@
where $w_{S\to r}$ $\in$ $\mathbb{R}^{d}$ is the identity latent and
@@
is the motion latent with $\lambda(S)$ :=:= $\left(\lambda_{m}(S)\right)_{m=1}^{M}$ $\in$ $\mathbb{R}^{M}$ being the source-dependent motion coefficients that span the learned source-agnostic motion basis $V:=\{\mathbf{v}_{m}\}_{m=1}^{M}\subseteq\mathbb{R}^{d}$. In this space, $\lambda_{m}(S)$ is the intensity of the motion direction $\mathbf{v}_{m}$. As shown in Fig.˜6, our method enables motion editing of the sampled (generated) motion using only the basis $V$ and its orthogonality, as stated in Eq.˜15.
@@ 4.2 Flow Matching in Motion Latent Space
Armed with this linear orthogonal space, we employ OT-based flow matching [42, 44] for the motion sampling. Specifically, we predict a vector field $v_{t}(x_{t},\mathbf{c}_{t};\theta)$ $\in$ $\mathbb{R}^{L\times d}$ where $x_{t}$ is the sample at flow time $t\in[0,1]$, and $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{L\times h}$ represents the driving conditions for $L$ consequent frames. This vector field generates the flow $\varphi_{t}:[0,1]\times\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ of $L$ frames by solving ODE (Eq.˜2). As illustrated in Fig.˜4, we build our vector field predictor upon the transformer encoder [79] architecture. Specifically, we adopt DiT [55] architecture, but decouple frame-wise conditioning from time-axis attention mechanism, which enables us to model temporally consistent motion latents.
@@
respectively, where $i$ $\in$ $\{1,2\}$, $h$ is the hidden dimension, $\text{LN}(\cdot)$ denotes layer norm [40], and $X_{t}^{l}$ is the $l$-th input for each operation at flow time $t$ $\in$ $[0,1]$. The coefficients $\alpha_{i}^{l},\beta_{i}^{l},\gamma_{i}^{l}\in\mathbb{R}^{h}$ are computed from the condition $\mathbf{c}_{t}^{l}$ $\in$ $\mathbb{R}^{h}$ through a linear layer, ToScaleShift, as depicted in Fig.˜4.
@@
Driving Conditions   We concatenate the audio representation $a^{1:L}$ $\in$ $\mathbb{R}^{L\times d_{a}}$ of a pre-trained Wav2Vec2.0 [1], the speech emotion label $w_{e}$ $\in$ $\mathbb{R}^{7}$, and the source motion latent $w_{r\to S}$ $\in$ $\mathbb{R}^{d}$. Next, we add the flow time step embedding $\text{Emb}(t)$ $\in$ $\mathbb{R}^{h}$ to these conditions, producing $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{L\times h}$ via a linear layer, ToCondition, as depicted in Fig.˜2, where $\text{Emb}(t)$ is computed using the sinusoidal position embedding [79].
Training   We train FLOAT by reconstructing a target vector field computed from driving frames using the corresponding audio segments and a source motion latent. We choose a pair of driving motions and corresponding audio ($w_{r\to D^{1:L}}$, $a^{1:L}$), and construct the target vector field $u_{t}(x|w_{r\to D^{1:L}})=w_{r\to D^{1:L}}-x_{0}$ $\in$ $\mathbb{R}^{L\times d}$ with noisy input $\varphi_{t}(x_{0})=(1-t)x_{0}+tw_{r\to D^{1:L}}$ ($t$ $\sim$ $\mathcal{U}[0,1]$ and $x_{0}$ $\sim$ $\mathcal{N}(0^{1:L},I)$).
@@
where $x_{t}$ :=:= $w_{r\to D^{-L^{\prime}:0}}|~\varphi_{t}(x_{0})] \in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$ is the concatenated input, $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times h}$ is the driving condition consisting of $t,w_{r\to S},w_{e},a^{1:L},a^{-L^{\prime}:0}]$. Note that $w_{e}$ and $w_{r\to S}$ are shared across the $L^{\prime}+L$ frames. We incorporate a velocity loss [75] to supervise temporal consistency:
@@
where $\Delta v_{t}$ and $\Delta u_{t}$ are the one-frame difference along the time-axis for the prediction $v_{t}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$ and the target $w_{r\to D^{-L^{\prime}:0}}|~u_{t}] \in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$, respectively.
@@
where $\lambda_{\text{OT}}$ and $\lambda_{\text{vel}}$ are the balancing coefficients. During training, we apply dropout to $w_{r}$, $w_{e}$, and $a^{1:L}$ with a probability of $0.1$ for CFV. Additionally, we apply dropout to the preceding audio and motion latents with a probability $0.5$ for smooth transition in the initial window.
Table 1: Quantitative comparison results with state-of-the-art methods on HDTF [97] / RAVDESS [46]. The best result for each metric is in bold, and the second-best result is underlined. : evaluated with raw $256\times 256$ resolution outputs.
@@ 5.2 Implementation Details
The motion latent dimension is set to $d$ == $512$ with $M$ == 2020 distinct orthogonal directions. For the vector predictor, we use 8 attention heads, a hidden dimension $h$ == $1024$, and an attention window length $T$ == 22. Considering the length of the training video clips, we set $L$ == $50$ frames with preceding $L^{\prime}$ == $10$ frames at once, encompassing 2.4 seconds of video. We employ the Adam optimizer [36] with a batch size of 8 and a learning late of $10^{-5}$. We use $L1$ distance for the norm $\|\cdot\|$ in the training objective. We set the balancing coefficients to $\lambda_{\text{OT}}$ == $\lambda_{\text{vel}}$ == 11. The entire training takes about $2$ days for 2,0002,000k steps on a single NVIDIA A100 GPU. We use Euler method [42] for the ODE solver.
@@ 5.4 Applications
Test-time Pose Editing via Orthonormal Basis $V$   Since FLOAT learns the underlying motion latent structure, it is natural to assume that for any sampled motion latent $w_{r\to\hat{D}}$, there exist motion coefficients $\{\lambda_{m}(\hat{D})\}_{m=1}^{M}$ satisfying the representation in Eq.˜9: $w_{r\to\hat{D}}$ == $\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m}$.
@@
[…]
@@
Additional Driving Signals   In Fig.˜7 and Tab.˜2, we experiment with additional driving conditions, head poses and image-driven emotion labels, to explore additional controllability in our method. We employ 3DMM head pose parameters $p$ $\in$ $\mathbb{R}^{6}$ [2] extracted by [13]. We concatenate a sequence of pose parameters $p^{1:L}$ $\in$ $\mathbb{R}^{L\times 6}$ with the other driving conditions, and then map them to $c_{t}^{1:L}$ $\in$ $\mathbb{R}^{L\times h}$. We also experiment on image-driven emotion [63] for frame-wise emotion control rather than the long-term emotion enhancement. FLOAT can effectively accommodate these additional conditions, highlighting its flexibility across diverse control signals.
@@ A.1 Model
Table 4: Quantitative comparison result (Same-identity) of motion latent auto-encoders on HDTF [97] / RAVDESS [46] / VFHQ [88]. The best result for each metric is in bold. : Results generated by official implementation ($256\times 256$)
@@ A.3 Training Objective
[…]
@@ B.2 Baselines
  • It is worth noting that we compare with two superior works EMO [76] and VASA-1 [90] based on their demonstration videos due to their unavailable implementation. We highly recommend referring to ‘01_\_EMO_\_VASA-1_\_Comparison//xxxx.mp4’.
@@ B.3 More on Experiments
Ablation on NFE   In general, increasing the number of function evaluation (NFE) reduces the solution error of ODEs. As shown in Tab.˜5, even with small NFE =2=2, FLOAT can achieve competitive image quality (FID) and lip synchronization (LSE-D). However, it struggles to capture consistent and expressive motions (FVD and E-FID), resulting in shaky head motion and a static expression. This is because FLOAT generates the motion in the latent space, while image fidelity is determined by the auto-encoder. We provide supplementary videos, illustrating the impact of different NFE (Number of Function Evaluations). Notably, with a small NFE of 2, the generated images exhibit good quality, but the head movements appear temporally unstable, and emotions may be exaggerated. Please refer to supplementary videos for temporal jitters of low NFE.
@@
In Fig.˜20, we visualize the effect of different emotion guidance scale $\gamma_{e}$. For this experiments, the predicted speech-to-emotion label is disgust with $99\%$ probability. Notably, as increasing $\gamma_{e}$ from 0 to 2, we can observe that emotion-related expressions and motions are enhanced.
@@
Ablation on AdaLN and Flow Matching  We conduct ablation study on frame-wise AdaLN by comparing it with a cross-attention. We adopt the stand cross-attention mechanism described in [19, 71], using transformer encoder architecture for non-autoregressive sequence modeling. We use the same attention mask used in the frame-wise AdaLN, which attends to additional $2T$ adjacent frames for the $l$-th input latent: [l2,l1,l,l+1,l+2][l-2,l-1,l,l+1,l+2].
@@
where $t$ $\sim$ $\mathcal{U}[0,1]$, $\epsilon$ $\sim$ $\mathcal{N}(0^{-L^{\prime}:L},I)$, and the noise input $x_{t}$ $\in$ $\mathbb{R}^{(L^{\prime}+L)\times d}$ is sampled from a forward diffusion process $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ [27]. In our case, $x_{t}$ is noisy motion latents at diffusion time step $t$, starting from $t=0$ with $x_{0}$ == $w_{r\to D^{1:L}}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$.